1)Deimos, a satellite of Mars, has an average radius of 6.2 km. If the

gravitational force between Deimos and a 2.0 kg rock at its surface is 2.5 ´10-2 N, what is the mass of Deimos?

 2. A 2.08 ´ 104 kg meteorite is on exhibit in New York City. Suppose this meteorite and another meteorite are separated by 1.27 ´ 107m (a distance equal to Earth’s average diameter). If the gravitational force between them is 2.88 ´ 10-16 N, what is the mass of the second meteorite?

 4. In 1989, a cake with a mass of 5.81 ´ 104 kg was baked in Alabama. Suppose a cook stood 25.0 m from the cake. The gravitational force exerted between the cook and the cake was 5.0 ´ 10-7 N. What was the cook’s mass?

 5. The largest diamond ever found has a mass of 621 g. If the force of gravitational attraction between this diamond and a person with a mass of

65.0 kg is 1.0 ´ 10-12 N, what is the distance between them?



1
Expert's answer
2022-01-14T09:46:39-0500

1.

"M=\frac{Fr^2}{Gm}=7.2\cdot 10^{15}~kg."

2.

"m_2=\frac{Fd^2}{Gm_1^2}=3.35\cdot 10^4~kg."

4.

"M=\frac{Fr^2}{Gm}=80.6~kg."

5.

"r=\sqrt{\frac{Gm_1m_2}F}=52~m."